2.. M. Chain was to manufacture 1,000 chain saws next month. Its accountant has provided the following analysis of the total manufacturing costs.

Variable Coefficient Standard Error t-Value
Constant 200 143.88 1.39
Predictor Variable 400 183.49 2.18

r2 = 0.71

What is the estimated cost of producing the 1,000 chain saws?

Select one:

A. $18,000

B. $284,142

C. $400,200

D. $9,000

E. $200,400

3.The demand for Ballard's Glass Company's products varies, ranging from 10 to 20 windows a day with an average of 15. John Ballard is the only worker and he works a maximum of eight hours a day, five days a week. Each order is one window, and each window takes 26 minutes to install. The company plans to add doors to its product line and anticipates selling an average of 5 doors per day. Each door takes 12 minutes to install. What is the average waiting time if Ballard is the only worker?

Select one:

A. 410.0 minutes

B. 38.0 minutes

C. 181.0 minutes

D. 112.4 minutes

E. 523.4 minutes3.

Hello,

Are you able to show work for both questions?

Thanks,
Answer #1

1. $400,200 is the estimated cost of producing the 1000 chain saws, calculations mentioned below:

Estimated cost = 200+400*1000

= 400,200

2. Waiting time = (15 * (26)^2) + (5+ (12)^2) / {2 * [480(minutes per day) - (15 * 26) - (5 * 12)]}

= 10860/60

= 181 minutes.

1. Option C is the correct answer which is $400,200.

2. Option C is the correct answer which is 181 minutes.